three objects A,B and C are at temperatures at 80C, 90C and 50C respectively. the objects are kept together in a calorimeter. assuming that heat is exchanged only among the three objects , find their final temperature.

Grade:10

2 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

The final temp will be average of 3 temps=80+90+50/3=220/3 C

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

gOlU g3n|[0]uS
13 years ago

ans... is the ans is 73.3o